# SUREL+: Moving from Walks to Sets for Scalable Subgraph-based Graph Representation Learning

### 17 Dec 2023 v2

@@ Appendix A Notations
Appendix B More Details
B.1. Other Related Works
Scalable GNN Design. GNNs are the most widely used tools for graph representation learning nowadays, although they face certain challenges when directly applied to subgraph-based methods. To address the scalability of GNNs, current studies focus on improving graph subsampling and mini-batch training techniques (Chiang et al., 2019; Zeng et al., 2020). However, graph subsampling used in GNNs fundamentally differs from subgraph extractions in SGRL. The goal of subsampling is to handle GPU memory overflow during full-batch training of GNN models. For SGRL, subgraphs sampled around a query serve as features for making predictions. Consequently, the scaling techniques developed for GNNs cannot be directly applied to SGRL. Another direction is to deploy distributed GNN systems for industry-level graphs. Unfortunately, these specialized techniques, including pipelining (Wan et al., 2022b), partitioned parallelism (Wan et al., 2022a), and update with staleness (Peng et al., 2022) do not address the main bottleneck of subgraph extraction for SGRL methods.
B.2. Model Design
Benefits of Subgraph-based Graph Representation Learning First, subgraph-based representation is versatile for different types of tasks, especially when queries of certain tasks involving multiple nodes and relations, e.g. existence of a link, property of a motif, development of higher-order patterns; while canonical GNNs are limited to handle such polyadic dynamics via node-wise representations (Srinivasan and Ribeiro, 2020; Wang and Zhang, 2021) (also refers to the example in Fig. 2). Second, subgraph-based models are more expressive by pairing with structural features to obtain most expressive structural node representations (Srinivasan and Ribeiro, 2020; Li et al., 2020; Wang and Zhang, 2021; Bouritsas et al., 2022). However, canonical GNNs cannot capture intra-distance information and joint relations over multiple nodes, which are critical to distinguishing nodes in structural symmetry and making predictions over them. Lastly, subgraph-based methods decouple the model depth from the receptive field since extracted subgraphs are localized to certain hops: when adding more layers for non-linearity, it does not contaminate embedding with irrelevant nodes or get over-smooth as canonical GNNs do. This results in a more robust representation and is particularly beneficial for modeling relations beyond singleton.
@@ B.3. Datasets
The full statistics of benchmark datasets are summarized in Table 7. OGB datasets [^1] https://ogb.stanford.edu/docs/dataset_overview/ are selected to benchmark our proposed framework and other baselines. It contains large-scale graphs (millions of nodes/edges) for real-world applications (i.e., academic and biological networks) and provides standard, open-sourced evaluation metrics and tools. Note that, vessel is a newly added benchmark of a biological graph, with $>3M$ nodes and sparse vessel structures extracted from the whole mouse brain (Paetzold et al., 2021), where nodes represent bifurcation points, and edges represent the blood vessels. Each node is associated with features of its physical location in the coordinate space $(x,y,z)$. The introduction of vessel provides a unique opportunity to examine graph representation learning (GRL) approaches in neuroscience, especially in scaling subgraph-based GRL methods to handle sparse and spatial graphs with millions of nodes and edges for scientific discovery.
criteo-click contains a sample of 30 days of Criteo live traffic data, each corresponding to one impression (a banner) displayed to a user and whether it is clicked (Diemert et al., 2017). Each record has 9 contextual features that are aggregated into a 270-dimensional edge feature. There are 675 unique campaign banners and 6.1M users, consisting of a bipartite graph of 16.5M edges: 97% is used for training, and the rest is evenly split for validation and testing based on temporal orders. The task is to predict which campaign the user is most likely to click among 651 candidates. twitter-2010 is an industry-level social network with 1.5B following relations of users (Kwak et al., 2010). An edge $(i,j)$ of this network indicates that user $i$ is followed by user $j$. 1% of Twitter users who follow 10 to 1000 accounts are randomly sampled for evaluation. The task is to recommend which account they will most likely follow among 1001 candidates. The OGB formatted files of these two datasets are accessible via Box at https://purdue.box.com/v/SGRL-LSC-dataset.
@@ B.4. Baselines
  • Canonical GCNN: a graph auto-encoder model that uses graph convolution layers to learn node representations, including GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017a), and their more scalable variants by employing graph subsampling, such as Cluster-GCN (Chiang et al., 2019), GraphSAINT (Zeng et al., 2020).
@@
  • SEAL [^4] https://github.com/facebookresearch/SEAL_OGB (Zhang and Chen, 2018): apply GCN on exact query-induced subgraphs with double radius node labeling to obtain subgraph-level readout for link prediction. SEAL shows great empirical performance on multiple graph machine learning benchmarks and promotes the deployment of subgraph-based models for scientific discovery. The implementation we tested is specialized for OGB datasets provided in (Zhang et al., 2021).
@@
  • GDGNN [^5] https://github.com/woodcutter1998/gdgnn (Kong et al., 2022): a model aggregates node representations generated by GNNs along geodesic paths between nodes in a query to speedup subgraph-based methods.
@@
  • SUREL [^6] https://github.com/Graph-COM/SUREL(Yin et al., 2022): a walk-based computation framework to accelerate subgraph-based methods, where subgraphs are decomposed to pre-sampled walks and then are joined online to substitute the query-induced subgraph for prediction. SUREL achieves state-of-the-art scalability and prediction accuracy on SGRL tasks by adopting walk-based representation.
@@
All canonical GNN baselines [^7] https://github.com/snap-stanford/ogb/tree/master/examples/linkproppred come with three GCNConv/SAGEConv layers of 256 hidden dimensions, and a tuned dropout ratio in $\{0,0.5\}$ for full-batch training. Canonical GNN models aggregate all node embeddings involved in a query as the representations of links/hyperedges, which are later fed into an MLP classifier for final prediction. In addition, all GNN models need to use full training data (edges/triplets) to generate robust node representations. The hypergraph datasets do not come with raw node features, thus GNN baselines here use randomly initialized features as input for training along with other model parameters. R-GCN uses RGCNConv layers that support message passing with multiple relation types between different types of nodes, where the edge types (relations) are used as input beside node features.
@@
Subgraph-based models only use partial edges/triplets for training. For SEAL, 1-hop enclosing subgraphs are extracted online during the training and inference. Then, it applies three GCN layers of 32 hidden dimensions plus a sort pooling and several 1D convolution layers to generate a readout of the target subgraph for prediction. SUREL consists of a 2-layer MLP for relative position encoding (RPE) and a 2-layer RNN to encode joined walks attached with RPEs to obtain the readout of query-level subgraphs. The hidden dimension of both networks is set to 64. The obtained embeddings of joined walks are aggregated and fed into a 2-layer MLP classifier to make predictions. GDGNN employs GINLayer as its backbone to obtain node embeddings. The horizontal geodesic representation is used for predictions, which finds the shortest path between two nodes in a query and aggregates node representations generated by GNNs along the found geodesic path. The max search distance for geodesic is the same as the number of GNN layers. For collab, ppa, citation2 and vessel, the threshold of distance is set to 4, 4, 3, and 2, respectively. The hidden dimension of all fully connected layers is set to 32.
@@
Appendix C Architecture and Hyperparameter
SUREL+ uses a 2-layer MLP with ReLU activation for encoding structural features and supports three set neural encoders, including mean pooling, LSTM, and attention. LSTM interprets elements to be aggregated in a set as a sequence (Hamilton et al., 2017a); attention first calculates soft attention scores for elements in a set and then performs attention-score-weighted average pooling. The hidden dimension of all parameterized layers is set to 96. Lastly, hidden representations of joined node sets are fed into a 2-layer MLP classifier for final predictions.
For the implementation of set samplers, the walk-based sampler builds on the sampling function from SubGAcc [^8] https://github.com/VeritasYin/subg_acc library developed by the authors, which also provides the support for efficient structural feature compression and index remapping. The metric-based sampler is adopted from fast PPR approximation in (Bojchevski et al., 2020).
We follow the inductive setting for link and relation prediction: only partial samples will be used for training. Over the training graph, we randomly select 5% links as positive training queries, each paired with $k$-many negative samples ($k=10$ by default). We mask these links and use the remaining 95% links to compute each node’s structural features in the split training set via structure encoder. For vessel, as the input graph is very sparse, we first sort the nodes in training set by their degree and then randomly pick 5% nodes to obtain edges of their 2-hop induced subgraphs for training and the rest reserved for structural feature construction. For higher-order pattern prediction, we use the given graph before timestamp $t$ to sample node sets and encode their structural features. The model parameters are optimized by triplets provided in the training set. No node features are used in SUREL+, except for vessel where normalized physical locations of each node are attached after its structural features and contextual features of click are attached.
The extended version of Table 4 is provided in Table 8 for the comparison on the rest of the datasets. The results reported in Table 3 and the profiling of SUREL+ in Tables 4, 8 are obtained through the combination of hyperparameters listed in Table 9. The dropout rate on vessel is set to p=0.2. The metric-based sampler is adopted to obtain the results of using PPR and SPD as structural features in Table 5. Its sampling size $K$ is set to $50$, $50$ and $150$ for collab, ppa, citation2, respectively. The walk-based sampler is used for the results of LP as structural features, whose sampling parameters are listed in Table 9. The rest of the hyperparameters remain the same as reported in Sec. 4.1. The SUREL+ framework including SubGAcc library is open-source and free for academic use under the BSD-2-Clause license.
b9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### 27 Dec 2023 v3

@@ B.3. Datasets
The full statistics of benchmark datasets are summarized in Table 7. OGB datasets [^1] https://ogb.stanford.edu/docs/dataset_overview/ are selected to benchmark our proposed framework and other baselines. The benchmark contains large-scale graphs (millions of nodes/edges) for real-world applications (e.g., academic and biological networks) and provides standard, open-sourced evaluation metrics and toolkits. Note that, vessel is a newly added benchmark of a biological graph, with $>3M$ nodes and sparse vessel structures extracted from the whole mouse brain (Paetzold et al., 2021), where nodes represent bifurcation points, and edges represent the blood vessels. Each node is associated with features of its physical location in the coordinate space $(x,y,z)$. The introduction of vessel provides a unique opportunity to examine graph representation learning approaches in neuroscience, especially in scaling subgraph-based methods to handle sparse and spatial graphs with millions of nodes and edges for scientific discovery.
criteo-click contains a sample of 30 days of Criteo live traffic data, each corresponding to one impression (a banner) displayed to a user and whether it is clicked (Diemert et al., 2017). Each record has 9 contextual features that are aggregated into a 270-dimensional edge feature. There are 675 unique campaign banners and 6.1M users, consisting of a bipartite graph of 16.5M edges: 97% is used for training, and the rest is evenly split for validation and testing based on temporal orders. The task is to predict which campaign the user is most likely to click among 651 candidates. twitter-2010 is an industry-level social network with 1.5B user following relations (Kwak et al., 2010). An edge $(i,j)$ of this network indicates that user $i$ is followed by user $j$. 1% of Twitter users who follow 10 to 1000 accounts are randomly sampled for evaluation. The task is to recommend which account they will most likely follow among 1001 candidates. The OGB formatted files of these two datasets are accessible via Box at https://purdue.box.com/v/SGRL-LSC-dataset.
@@ B.4. Baselines
  • Canonical GNNs: a graph auto-encoder model that uses graph convolution layers to learn node-wise representations, including GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017a), and their more scalable variants by employing graph subsampling, such as GraphSAINT (Zeng et al., 2020).
@@
  • SEAL [^4] https://github.com/facebookresearch/SEAL_OGB (Zhang and Chen, 2018): apply GCN on query-induced subgraphs attached with double radius node labeling to obtain subgraph-level readout for link prediction. SEAL shows great empirical performance on multiple graph machine learning benchmarks and promotes the deployment of subgraph-based models for scientific discovery. The implementation we tested is specialized for OGB datasets provided in (Zhang et al., 2021).
@@
  • GDGNN [^5] https://github.com/woodcutter1998/gdgnn (Kong et al., 2022): a subgraph-based model aggregates node representations generated by GNNs along geodesic paths between queried nodes for fast inference.
@@
  • SUREL [^6] https://github.com/Graph-COM/SUREL(Yin et al., 2022): a walk-based computation framework to accelerate subgraph-based methods, where subgraphs are decomposed to pre-sampled walks and then are joined online to substitute the query-induced subgraph for prediction. By adopting the walk-based representation, SUREL achieves state-of-the-art scalability and prediction accuracy on SGRL tasks.
@@
All canonical GNN baselines [^7] https://github.com/snap-stanford/ogb/tree/master/examples/linkproppred come with three GCNConv/SAGEConv layers of 256 hidden dimensions, and a tuned dropout ratio in $\{0,0.5\}$ for full-batch training. Canonical GNNs aggregate all node embeddings involved in a query as the representation of link/hyperedge, which is later fed into an MLP classifier for final prediction. In addition, all GNN models need to use full training data (edges/triplets) to generate robust node representations. The hypergraph datasets do not come with raw node features, and thus GNN baselines use randomly initialized features as input for training along with other model parameters. R-GCN uses RGCNConv layers that support message passing with multiple relation types between different types of nodes, where the edge types (relations) are used as input besides node features.
@@
Subgraph-based models only use partial edges/triplets for training. For SEAL, 1-hop enclosing subgraphs are extracted online during the training and inference. Then, it applies three GCN layers of 32 hidden dimensions plus a sort pooling and several 1D convolution layers to generate a readout of the target subgraph for prediction. SUREL consists of a 2-layer MLP for query-level relative position encoding (RPE) and a 2-layer RNN to encode joined walks with attached RPEs. The hidden dimension of both networks is set to 64. The obtained readout of joined walks is aggregated and fed into a 2-layer MLP classifier to make predictions. GDGNN employs GINLayer as its backbone to obtain node embeddings. The horizontal geodesic representation is used for predictions, which finds the shortest path between two nodes in a query and aggregates node representations generated by GNNs along the found geodesic path. The max search distance for geodesic is the same as the number of GNN layers. For collab, ppa, citation2 and vessel, the threshold of distance is set to 4, 4, 3, and 2, respectively. The hidden dimension of all fully connected layers is set to 32.
Appendix C Architecture and Hyperparameter
SUREL+ uses a 2-layer MLP with ReLU activation for encoding structural features and supports three set neural encoders, including mean pooling, LSTM, and attention. LSTM interprets elements to be aggregated in a set as a sequence (Hamilton et al., 2017a); attention first calculates soft attention scores for elements in a set and then performs attention-score-weighted average pooling. The hidden dimension of all parameterized layers is set to 96. Lastly, hidden representations of query-level joined node sets are fed into a 2-layer MLP classifier for final predictions.
The walk-based sampler builds on the sampling function from SubGAcc [^8] https://github.com/VeritasYin/subg_acc library developed by the authors, which also provides the support for efficient structural feature compression and index remapping. The metric-based sampler is adopted from fast PPR approximation in (Bojchevski et al., 2020).
@@ Appendix C Architecture and Hyperparameter
We follow the inductive setting for link and relation prediction: only partial samples will be used for training. Over the training graph, we randomly select 5% links as positive training queries, each paired with $k$-many negative samples ($k=10$ by default). We mask these links and use the remaining 95% links to compute each node’s structural features in the split training set via structure encoder. For vessel, as the input graph is very sparse, we first sort the nodes in training set by their degree and then randomly pick 5% nodes to obtain edges of their 2-hop induced subgraphs for training and the rest reserved for structural feature construction. For higher-order pattern prediction, we use the given graph before timestamp $t$ to sample node sets and encode their structural features. The model parameters are optimized by triplets provided in the training set. No node features are used in SUREL+, except for vessel where normalized physical locations of each node are attached after its structural features and similarly for contextual features in click.
Table 8 presents the extended version of Table 4. The results reported in Table 3 and the profiling of SUREL+ in Tables 4, 8 are obtained through the combination of hyperparameters listed in Table 9. The dropout rate on vessel is set to p=0.2. The metric-based sampler is adopted to obtain the results of using PPR and SPD as structural features in Table 5. Its sampling size $K$ is set to $50$, $50$ and $150$ for collab, ppa, citation2, respectively. The walk-based sampler is used for the results of LP as structural features, whose sampling parameters are listed in Table 9. The rest of the hyperparameters remain the same as reported in Sec. 4.1. The SUREL+ framework including SubGAcc library is open-source and free for academic use under the BSD-2-Clause license.